A $15000 investment is to be divided among two stocks so that $3000 less is invested in one stock than the other. How much will be invested in each.
I have to make an equation for this problem. I'm a little stumped.

Answer by edjones(8007)

x is invested in one and x-3000
x+x-3000=15000
2x-3000=15000
2x=18000
x=$9000
x-3000=$6000
.
Ed